### v1

# LIBERO-PRO: Towards Robust and Fair Evaluation of Vision-Language-Action Models Beyond Memorization

Abstract
LIBERO has emerged as a widely adopted benchmark for evaluating Vision-Language-Action (VLA) models; however, its current training and evaluation settings are problematic, often leading to inflated performance estimates and preventing fair model comparison. To address these issues, we introduce LIBERO-PRO, an extended LIBERO benchmark that systematically evaluates model performance under reasonable perturbations across four dimensions: manipulated objects, initial states, task instructions, and environments. Experimental results reveal that, although existing models achieve over 90% accuracy under the standard LIBERO evaluation, their performance collapses to 0.0% under our generalized setting. Crucially, this discrepancy exposes the models’ reliance on rote memorization of action sequences and environment layouts from the training set, rather than genuine task understanding or environmental perception. For instance, models persist in executing grasping actions when the target object is replaced with irrelevant items, and their outputs remain unchanged even when given corrupted instructions or even messy tokens. These findings expose the severe flaws in current evaluation practices, and we call on the community to abandon misleading methodologies in favor of robust assessments of model generalization and comprehension. Our code is available at: https://github.com/Zxy-MLlab/LIBERO-PRO.
Figure 1: Model trajectory consistency under object and instruction modifications. Original task” denotes the standard evaluation, while “Replace,” “Change Pos.,” “Remove,” and “Messy instruction” introduce perturbations to the target object or instruction. Across all settings, the model yields nearly identical trajectories, suggesting a lack of genuine task understanding and environmental perception.
1 INTRODUCTION
Vision-Language-Action (VLA) models have recently emerged as a central paradigm for bridging perception, language understanding and action execution in embodied agentsSapkota et al. (2025). A crucial factor in advancing this line of research is the availability of standardized benchmarksJames et al. (2020); Liu et al. (2023); Mees et al. (2022); Li et al. (2024b); Nasiriany et al. (2024); Walke et al. (2023) that allow fair and reproducible comparison across different methods. Among them, LIBEROLiu et al. (2023) has rapidly become the most widely adopted evaluation suite for VLA, serving as the de facto standard for reporting performance. As shown in Table 1, models ranging from large-scale pre-trained to specialized systems are all benchmarked on LIBERO, making it the “common currency” of evaluation in this field. Consequently, the reliability of LIBERO as an evaluation protocol plays a decisive role in shaping research directions in VLA.
Figure 2: Success rates of OpenVLA, pi0, pi0.5, and univla under object position perturbations.
However, the current LIBERO training–evaluation setup may suffer from serious limitations. In fact, the evaluation tasks are exactly the same as the training tasks, with the only difference being slight perturbations in the initial states of objects—perturbations that are almost invisible to the human eye. As a result, if a model overfits the training set, it can easily achieve high scores during evaluation, but its accuracy drops to nearly zero under minor changes in object initial positions or task instructions, as shown in Figure 2. Yet such results do not indicate that the model has learned to understand the environment or the task. On the contrary, we find that models rely on fixed mappings from the training tasks. As illustrated in Figure 1, when we introduced four types of modifications—(1) replacing the target object with an unrelated one; (2) relocating its initial position; (3) removing it entirely; and (4) corrupting the instruction with nonsensical tokens—their trajectories remained nearly identical. This outcome shows that, despite achieving over 90% on the standard LIBERO benchmark, the reported scores reflect rote recall rather than genuine task understanding. Obviously, the current evaluation protocol is inadequate and fails to effectively assess model architectures and algorithms.
To address the above shortcomings, we systematically propose an extended evaluation suite, i.e., LIBERO-PRO, to enable a more comprehensive and reasonable assessment of VLA models. We design evaluation scenarios along four key dimensions: manipulated objects, initial states, task instructions, and environments. While ensuring that each task remains coherent and executable, we introduce carefully designed perturbations to the original LIBERO tasks to simulate real-world situations. Using this suite, we evaluate state-of-the-art models including OpenVLAKim et al. (2024) and pi0/0.5, and the results show that they exhibit near-complete failure under several of these settings: the models cannot cope with changes in object positions, nor can they correctly follow slightly rephrased task instructions—even when such instructions have appeared in other training contexts. These findings reveal that the commonly reported accuracies of over 90% are largely artifacts of task memorization, rather than evidence of genuine perception or task understanding. We therefore call on the community to adopt evaluation standards that incorporate such necessary perturbations, ensuring that reported progress reflects capabilities required for practical use. Our main contributions are as follows:
  • Through extensive experiments, we demonstrate that the current LIBERO evaluation protocol is fundamentally flawed: high scores primarily reflect rote memorization of training data rather than genuine task understanding or execution ability.
  • We introduce LIBERO-PRO, a plug-and-play benchmark that incorporates carefully designed perturbations across four dimensions—manipulated objects, initial states, task instructions, and environments—and supports randomized combinations of these dimensions for more faithful evaluation.
  • We benchmark state-of-the-art models, including OpenVLA, pi0 and pi0.5, on LIBERO-PRO and show that they perform poorly, highlighting the inadequacy of the current LIBERO protocol and calling for LIBERO-PRO as a fairer and more reliable evaluation standard.
2 Related Work
Vision-Language-Action Models. Vision-Language-Action (VLA) models have rapidly emerged as a central paradigm for embodied AI, aiming to unify perception, natural language understanding, and action execution within a single framework. Recent years have witnessed an explosion of approaches, ranging from large-scale pre-trained models that leverage multimodal corpora to task-specific architectures tailored for robotic manipulation and navigationMa et al. (2025); Li et al. (2025b); Din et al. (2025). While these efforts have significantly advanced the field, they also underscore the urgent need for standardized evaluation benchmarksLiu et al. (2023); Mees et al. (2022); James et al. (2020). Without such benchmarks, it becomes difficult to conduct fair and reproducible comparisons across methods, hindering the cumulative progress of VLA research.
Benchmarks for VLA Evaluation. A number of benchmarks have been developed to enable systematic evaluation of VLA models, including RLBench James et al. (2020), CALVIN Mees et al. (2022), RoboCasa Nasiriany et al. (2024), and BridgeData Walke et al. (2023). These benchmarks provide valuable testbeds for task execution under vision and language guidance, yet their adoption has remained relatively limited within the broader community. In contrast, LIBERO Liu et al. (2023) has quickly emerged as the dominant benchmark and is now the most widely adopted protocol for evaluating VLA models. Its standardized task suite and unified reporting metrics have made it the de facto standard for performance comparison, to the extent that virtually all recent VLA studies report results on LIBERO. As such, LIBERO not only serves as the “common currency” of evaluation in this field, but also exerts a strong influence on research directions and claims of progress.
Table 1: Incomplete statistics of VLA models evaluated on LIBERO. Numbers shown to two decimal places. ‘–’ denotes not reported. Single: One checkpoint trained for all tasks. Multi: Separate checkpoint trained for each task.
Model Year Size Method Single/Multi Goal Object Spatial 10 Average
pi0.5 Intelligence et al. (2025) 2025 $\sim$7B SFT Single 0.98a 0.98 0.98a 0.92 0.97
CogACT Li et al. (2024a) 2024 $\sim$7B SFT - 0.90 0.98b 0.97b 0.89 0.98a
GeoVLA Sun et al. (2025) 2025 $\sim$7B SFT - 0.97b 0.99a 0.98a 0.97a 0.98a
OpenVLA Kim et al. (2024) 2024 $\sim$7B SFT Multi 0.98a 0.99a 0.98a 0.93b 0.97b
pi0 Black et al. (2024) 2024 $\sim$7B SFT Single 0.92 0.98b 0.97b 0.96b
UniVLA* Wang et al. (2025b) 2025 $\sim$8B SFT Single 0.94 0.99a 0.95 0.94b 0.96b
SwitchVLA Li et al. (2025c) 2025 SFT Single 0.93 0.93
TGRPO Chen et al. (2025) 2025 $\sim$7B RL Multi 0.91 0.91
GraspVLA Deng et al. (2025) 2025 $\sim$2B SFT Single 0.91 0.94 0.82 0.89
UniVLA Bu et al. (2025) 2025 $\sim$7B SFT Multi 0.90 0.94 0.91 0.79 0.89
TriVLA Liu et al. (2025) 2025 $\sim$3B SFT - 0.90 0.94 0.91 0.73 0.87
MolmoAct Lee et al. (2025) 2025 $\sim$7B SFT Multi 0.88 0.95 0.87 0.77 0.87
CronusVLA Li et al. (2025a) 2025 $\sim$7B SFT Multi 0.91 0.95 0.90 0.69 0.86
SmolVLA Shukor et al. (2025) 2025 $\sim$2B SFT Single 0.91 0.94 0.93 0.73 0.86
ThinkAct Huang et al. (2025) 2025 $\sim$7B SFT+RL - 0.87 0.91 0.88 0.71 0.84
World-VLA Cen et al. (2025) 2025 SFT Multi 0.83 0.96 0.88 0.60 0.82
Cot-VLA Zhao et al. (2025) 2025 $\sim$7B SFT - 0.88 0.92 0.88 0.69 0.81
VLA-RL Lu et al. (2025) 2025 $\sim$7B SFT+RL Multi 0.82 0.92 0.90 0.60 0.81
SpatialVLA Qu et al. (2025) 2025 $\sim$4B SFT Multi 0.79 0.90 0.88 0.56 0.78
Octo Team et al. (2024) 2024 $\sim$93M SFT Multi 0.85 0.86 0.79 0.51 0.75
VQ-VLA Wang et al. (2025c) 2025 $\sim$7B SFT Single 0.75 0.60 0.68
  • aBest (cell shaded light blue).   bSecond-best (cell shaded light green).
Limitations of Current Evaluation Practices. Several recent studies have noted that the high scores reported on LIBERO often fail to translate into reliable performance, even on tasks with only minor variations. This has raised serious concerns about the validity of its evaluation protocolZhou et al. (2025); Fang et al. (2025); Li (2025). In response, some researchers have explored stronger algorithms or proposed stress tests Gu et al. (2025); Guruprasad et al. (2024); Valle et al. (2025); Kube et al. (2025); Wang et al. (2025a)to encourage more robust VLA systems. However, the field still lacks a systematic and standardized framework for robust evaluation under LIBERO. The majority of works continue to rely on the original, flawed setup, reporting numbers that may not reflect genuine capability. To address this gap, we propose LIBERO-PRO, a plug-and-play extension of LIBERO designed to foster fair comparison across models and to advance research on building VLA systems that can withstand reasonable and practically necessary perturbations.
3 Problem Definition
This section formalizes the Vision-Language-Action (VLA) task framework and examines limitations in the LIBERO benchmark’s evaluation protocol that may obscure true task competence.
3.1 Formal Task Definition
We define a VLA task to establish a rigorous basis for evaluating robotic systems.
$\tau=(\ell,\mathcal{O},\mathcal{E},p_{0},G),$ (1)
where $\ell$ is a natural language instruction specifying the task objective, $\mathcal{O}$ is a finite set of objects with attributes (e.g., shape, color, size, initial pose), $\mathcal{E}$ denotes the static environmental context (e.g., scene geometry, furniture layout), $p_{0}$ is a probability distribution over initial states $s_{0}$ (encoding init position and robot proprioception), and $G(s)$ is a binary goal predicate, returning $1$ if state $s$ satisfies $\ell$, and 0 otherwise.
The performance of a policy $\pi_{\theta}(a\mid s,\ell)$, parameterized by $\theta$, is measured by the expected success probability over a finite horizon $T$:
$\mathbb{P}_{s_{0}\sim p_{0},\pi_{\theta}}\left[G(s_{T})=1\right],$ (2)
where $s_{T}$ is the terminal state after $T$ steps. This formulation captures the integration of perception, language understanding, and action planning in VLA systems.
3.2 Current LIBERO Evaluation Protocol
The LIBERO benchmark, while valuable for VLA research, employs an evaluation protocol that may overestimate model competence due to its homogeneous train/test split. Each test task $\tau^{\prime}\in\mathcal{T}_{\mathrm{test}}$ mirrors a training task $\tau\in\mathcal{T}_{\mathrm{train}}$:
$\ell^{\prime}=\ell,\quad\mathcal{O}^{\prime}=\mathcal{O},\quad\mathcal{E}^{\prime}=\mathcal{E},\quad G^{\prime}=G,$ (3)
with the only difference being a minor perturbation in the initial state distribution:
$d_{\mathrm{TV}}(p_{0},p_{0}^{\prime})\leq\epsilon,\quad\epsilon\ll 1.$ (4)
Due to the small $\epsilon$, test tasks closely resemble training tasks, allowing policies to achieve high performance through memorization rather than robust generalization. This homogeneity may inflate performance metrics, masking limitations in compositional generalization and adaptability to diverse real-world scenarios, a critical concern for embodied AI.
Figure 3: LIBERO-PRO benchmark task overview. We extend the four original task categories in LIBERO by introducing object attribute perturbations, initial position perturbations, instruction perturbations, and environmental perturbations.
4 Methodology
4.1 Proposed Perturbation-Based Evaluation Framework
For a more reasonable evaluation of VLA models, we propose a framework that introduces controlled perturbations across four orthogonal dimensions, manipulated objects, initial states, task instructions, and environments. For a task $\tau=(\ell,\mathcal{O},\mathcal{E},p_{0},G)$, we generate perturbed tasks $\tau^{(k)}=\phi_{k}(\tau)$ using operators $\phi_{k}$, $k\in\{\mathrm{O},\mathrm{S},\mathrm{L},\mathrm{E}\}$, defined as follows:
Object Attribute Perturbation ($k=\mathrm{O}$): Modifies non-essential object attributes, such as color, texture, or size, while preserving semantic equivalence, e.g., changing a red cup to a blue cup. This tests the model’s robustness to superficial visual changes critical for real-world adaptability. The perturbed task is:
$\tau^{(\mathrm{O})}=\left(\ell,\mathcal{O}^{(\mathrm{O})},\mathcal{E},p_{0},G\right).$ (5)
Initial Position Perturbation ($k=\mathrm{I}$): Alters initial object positions to modify absolute and relative arrangements, ensuring physical plausibility, e.g., moving a cup relative to a plate. This probes the model’s spatial reasoning under varied layouts. The perturbed task is:
$\tau^{(\mathrm{I})}=\left(\ell,\mathcal{O},\mathcal{E},p_{0}^{(\mathrm{I})},G\right).$ (6)
Instruction Perturbation ($k=\mathrm{L}$): Perturbations are applied at both the semantic and task levels. Semantic perturbations adjust the linguistic expression while preserving the original task intent, such as paraphrasing “pick up the bottle” to “grab the bottle” or substituting with a jar. This evaluates the agent’s understanding of language and semantics. The perturbed task is:
$\tau^{(\mathrm{L})}=\left(\ell^{(\mathrm{L})},\mathcal{O},\mathcal{E},p_{0},G\right),$ (7)
Task perturbations reasonably modify the task itself by replacing the target object or altering the action, e.g., changing “pick up the bottle” to “pick up the bowl.” We ensure that all actions and objects appear in the training set to support fair and meaningful generalization. The perturbed task is:
$\tau^{(\mathrm{L})}=\left(\ell^{(\mathrm{L})},\mathcal{O}^{(\mathrm{L})},\mathcal{E},p_{0},G^{(\mathrm{L})}\right).$ (8)
Environment Perturbation ($k=\mathrm{E}$): Modifies the background’s appearance, e.g., changing a workspace, without affecting task feasibility, to assess robustness to environmental variations. The perturbed task is:
$\tau^{(\mathrm{E})}=\left(\ell,\mathcal{O},\mathcal{E}^{(\mathrm{E})},p,G\right).$ (9)
4.2 Perturbation Constraints
To ensure each perturbed task $\tau^{(k)}$ remains semantically valid, we impose two constraints. First, perturbations are confined within a dimension-specific neighborhood:
$x^{(k)}\in\mathcal{N}_{k}(x),\quad\mathrm{dist}_{k}(x,x^{(k)})\leq\delta_{k},$ (10)
where $x$ is the original component, $x^{(k)}$ is its perturbed version, and $\delta_{k}$ controls the magnitude of perturbation for dimension $k\in\{\mathrm{O},\mathrm{S},\mathrm{L},\mathrm{E}\}$. Second, we enforce non-trivial task-level separation:
$d_{\mathrm{TV}}(\tau,\tau^{(k)})>\epsilon,\quad\forall k\in\{\mathrm{O},\mathrm{S},\mathrm{L},\mathrm{E}\},$ (11)
where $d_{\mathrm{TV}}$ is the total variation distance, ensuring that perturbations lead to substantive deviations. Multiple perturbed tasks are generated per category by applying $\phi_{k}$, with $\delta_{k}$ varied to characterize the sensitivity of policies under different perturbation magnitudes.
4.3 LIBERO-PRO Evaluation Benchmark
Building on these principles, we introduce LIBERO-PRO, a plug-and-play evaluation benchmark that enables systematic assessment under multiple dimensions of variation—object, position, instruction, task, and environment—as well as their arbitrary combinations, as shown in Figure 3. Specifically, for object variation, we create new assets for the four LIBERO tasks by modifying the appearance, size, and color of the original objects. For position variation, we define alternative spatial regions for all manipulable objects while strictly adhering to physical constraints and task definitions. For instruction variation, we provide three distinct paraphrased versions of each task description. For task variation, we redesign tasks in a principled manner to ensure feasibility, and supply new sets of relevant objects and target states for each modified task. For environment variation, we leverage the five environment categories built into the LIBERO benchmark and randomly substitute them across tasks. To avoid meaningless combinations, we restrict task variation from being combined with other types, while allowing all other dimensions of variation to be freely and randomly combined for more comprehensive evaluation.
5 Experiments
5.1 Settings
We conduct experiments on LIBERO-PRO, which applies controlled perturbations from four perspectives: object properties, initial positions, task instructions (covering both semantic and structural variations), and environments. We evaluate three representative models—OpenVLAKim et al. (2024), pi0Black et al. (2024), and pi0.5Intelligence et al. (2025). For fairness, OpenVLA is assessed using four official checkpoints trained on different LIBERO suites, whereas pi0 and pi0.5 are evaluated with a single official checkpoint. Consistent with the original LIBERO protocol, we set the number of evaluation episodes to 50 per task.
5.2 Main Results
We conducted a comprehensive evaluation of state-of-the-art vision-language-action (VLA) models Black et al. (2024); Kim et al. (2024) using the LIBERO-PRO evaluation suite, which introduces perturbations across four dimensions: objects, goals, spatial configurations, and task sets (see Tables 2, 3, 4, 5). Our findings are as follows: (1) Failure under simple perturbations. Despite achieving success rates above 90% on the standard LIBERO benchmark, models nearly collapse under changes to object positions or minor task modifications, even when constructed from training components. (2) Illusory robustness. Apparent stability against object and instruction perturbations is largely due to overfitting rather than genuine task understanding. (3) Model-specific sensitivity. Environmental perturbations affect models differently, depending on the degree of environmental divergence and reliance on memorized contexts. (4) Subtle capability differences. While most models fail under perturbations, Pi0.5 achieves a 0.38 success rate in the libero-goal task under position changes, compared to 0 for OpenVLA and Pi0, revealing differences masked by standard benchmark scores. Overall, these results show that current VLA models, despite excelling on the standard LIBERO, lack true robustness and generalization—highlighting the need for rigorous evaluation frameworks.
Table 2: Task-wise LIBERO and LIBERO-PRO performance of three VLA models on the libero-goal benchmark. Accuracy on test (Ori: origina, Obj: object, Spa: spatial, Sem: semantic, Task: task-level, Env: environment). Action notation: $Open(x,y)$ = open target $y$ of container $x$; $Put(obj,loc)$ = place object objobj onto/into location locloc; $Push(obj,loc)$ = push object objobj toward location locloc; $TurnOn(obj)$ = activate object objobj.
Task (Symbolic Form) OpenVLA Pi0 Pi0.5
Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env
$Open(cabinet,drawer_{mid})$ 0.98 0.96 0.00 1.00 0.00 0.96 0.92 0.94 0.00 0.94 0.00 0.00 0.96 0.96 0.00 0.94 0.04 0.92
$Put(bowl,drawer_{top})$ 0.88 0.88 0.00 0.96 0.00 0.82 0.76 0.86 0.00 0.94 0.00 0.10 0.98 0.98 0.94 1.00 0.02 1.00
$Push(plate,stove_{front})$ 1.00 1.00 0.00 1.00 0.00 1.00 0.94 0.93 0.00 0.94 0.00 0.54 0.96 0.98 0.00 0.98 0.00 0.28
$Put(bowl,plate)$ 1.00 1.00 0.00 1.00 0.00 1.00 1.00 0.99 0.00 0.96 0.00 0.96 0.90 0.90 0.00 0.92 0.02 0.04
$Put(bowl,stove)$ 0.94 1.00 0.00 1.00 0.00 1.00 0.98 0.97 0.00 0.98 0.00 0.98 0.98 0.96 0.00 0.98 0.04 0.78
$Put(bowl,cabinet_{top})$ 1.00 0.98 0.00 1.00 0.00 1.00 0.98 0.94 0.00 0.92 0.00 0.94 0.96 0.96 0.00 0.96 0.02 0.00
Put(cream_cheese,bowl)Put(cream\_cheese,bowl) 1.00 1.00 0.00 1.00 0.00 1.00 0.96 0.96 0.00 0.96 0.00 0.46 0.98 1.00 0.98 0.98 0.02 0.94
Put(wine_bottle,rack)Put(wine\_bottle,rack) 1.00 1.00 0.00 1.00 0.00 1.00 0.68 0.84 0.00 0.64 0.00 0.00 1.00 0.98 0.88 0.96 0.02 0.12
$Put(wine\_bottle,cabinet_{top})$ 0.96 0.83 0.00 0.88 0.00 0.88 0.94 1.00 0.00 1.00 0.00 0.00 1.00 1.00 0.98 0.98 0.02 0.48
$TurnOn(stove)$ 1.00 1.00 0.00 1.00 0.00 1.00 1.00 1.00 0.00 1.00 0.00 0.00 1.00 1.00 0.00 0.96 0.00 0.00
Average 0.98 0.96 0.00 0.98 0.00 0.98 0.92 0.94 0.00 0.93 0.00 0.39 0.97 0.97 0.38 0.97 0.00 0.46
Table 3: Task-wise LIBER0 and LIBERO-PRO performance of three VLA models on the libero-spatial benchmark. Action notation: $Pick(src)$ = pick up object $bowl_{black}$ from location srcsrc and place $bowl_{black}$ onto/into target dstdst.
Task (Symbolic Form) OpenVLA Pi0 Pi0.5
Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env
$Pick(between(plate,ramekin),plate)$ 1.00 0.95 0.00 0.88 0.00 0.02 1.00 0.97 0.00 0.94 0.00 1.00 1.00 0.96 0.02 0.98 0.00 0.66
Pick(table_center,plate)Pick(table\_center,plate) 1.00 0.98 0.00 1.00 0.00 1.00 1.00 1.00 0.00 0.98 0.00 0.98 1.00 1.00 0.00 0.98 0.02 0.36
$Pick(drawer_{top}(cabinet_{wood}),plate)$ 0.96 0.96 0.00 0.94 0.00 0.94 0.98 0.99 0.00 0.94 0.00 0.54 1.00 1.00 0.00 1.00 0.00 0.64
Pick(next_to(cookie_box),plate)Pick(next\_to(cookie\_box),plate) 1.00 1.00 0.00 1.00 0.00 1.00 1.00 0.94 0.00 1.00 0.00 1.00 1.00 0.98 0.00 1.00 0.02 0.86
Pick(next_to(plate),plate)Pick(next\_to(plate),plate) 1.00 0.95 0.00 0.94 0.00 0.92 0.96 1.00 0.00 0.94 0.00 0.84 0.90 0.92 0.00 0.90 0.00 0.22
Pick(next_to(ramekin),plate)Pick(next\_to(ramekin),plate) 1.00 0.99 0.00 1.00 0.00 0.98 1.00 0.80 0.00 1.00 0.00 0.48 1.00 0.96 0.12 1.00 0.02 0.00
Pick(on(cookie_box),plate)Pick(on(cookie\_box),plate) 1.00 0.99 0.00 0.98 0.00 1.00 0.98 0.92 0.00 0.98 0.00 0.96 1.00 1.00 0.00 0.98 0.00 1.00
$Pick(on(ramekin),plate)$ 0.92 0.95 0.00 0.96 0.00 1.00 0.90 1.00 0.00 0.96 0.00 0.10 1.00 1.00 0.98 1.00 0.02 0.00
$Pick(on(stove),plate)$ 1.00 0.99 0.00 1.00 0.00 1.00 0.96 0.99 0.00 0.96 0.00 0.04 0.96 0.96 0.02 0.94 0.00 0.82
$Pick(on(cabinet_{wood}),plate)$ 0.94 0.99 0.00 1.00 0.00 1.00 0.94 0.86 0.00 1.00 0.00 0.80 0.90 0.88 0.90 0.90 0.00 0.02
Average 0.98 0.97 0.00 0.97 0.00 0.89 0.97 0.95 0.00 0.97 0.00 0.60 0.98 0.97 0.20 0.97 0.01 0.46
Table 4: Task-wise LIBERO and LIBERO-PRO performance of three VLA models on the libero-10 benchmark. Action notation: $TurnOn(obj)$ = activate object objobj; $Put(obj,loc)$ = place object objobj onto/into location locloc; $Place(obj,loc)$ = place picked object objobj into/onto target locloc; $(obj).close$ = close container objobj.
Task (Symbolic Form) OpenVLA Pi0 Pi0.5
Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env
$TurnOn(stove)\wedge Put(moka\_pot,stove)$ 1.00 0.95 0.00 1.00 0.00 0.00 0.74 0.67 0.00 0.74 0.00 0.00 0.92 0.90 0.18 0.90 0.00 0.94
$Put(bowl_{black},drawer_{bottom}(cabinet)).close$ 0.78 0.98 0.00 1.00 0.00 0.98 0.94 0.93 0.00 0.94 0.00 0.78 0.98 0.96 0.58 0.98 0.02 0.86
$Put(mug_{yellow\_white},microwave).close$ 1.00 0.98 0.00 0.94 0.00 0.96 0.88 0.74 0.00 0.88 0.00 0.00 0.98 0.98 0.00 1.00 0.00 0.46
$Put(\{moka\_pot_{1},moka\_pot_{2}\},stove)$ 0.92 0.64 0.00 0.92 0.00 0.78 0.22 0.28 0.00 0.22 0.00 0.00 0.98 0.98 0.00 0.98 0.02 0.26
$Put(\{alphabet\_soup,cream\_cheese\},basket)$ 0.96 0.98 0.00 1.00 0.00 1.00 1.00 0.96 0.00 1.00 0.00 0.16 1.00 0.98 0.00 0.98 0.00 0.52
$Put(\{alphabet\_soup,tomato\_sauce\},basket)$ 0.98 0.95 0.00 0.98 0.00 0.96 0.88 0.79 0.00 0.88 0.00 0.06 0.94 0.94 0.02 0.94 0.02 0.70
$Put(\{cream\_cheese,butter\},basket)$ 0.96 0.96 0.00 0.98 0.00 0.98 0.98 0.98 0.00 0.98 0.00 0.96 0.96 0.94 0.00 0.98 0.00 0.00
$Put((mug_{white},plate_{left})|(mug_{yellow\_white},plate_{right}))$ 0.84 0.85 0.00 0.92 0.00 0.98 0.86 0.72 0.00 0.86 0.00 0.00 1.00 0.98 0.00 1.00 0.02 0.22
$Put((mug_{white},plate)|(pudding_{choco},right\_of(plate)))$ 0.94 0.77 0.00 0.90 0.00 0.86 0.76 0.82 0.00 0.74 0.00 0.00 0.62 0.58 0.00 0.60 0.00 0.04
$Place(book,compartment_{back}(caddy))$ 1.00 1.00 0.00 0.98 0.00 1.00 0.94 0.98 0.00 0.94 0.00 0.76 0.94 0.94 0.00 0.94 0.00 0.56
Average 0.93 0.81 0.00 0.96 0.00 0.85 0.82 0.79 0.00 0.82 0.00 0.27 0.93 0.92 0.08 0.93 0.01 0.46
Table 5: Task-wise LIBERO and LIBERO-PRO performance of three VLA models on the libero-object benchmark. Action notation: $Place(obj,loc)$ = pick up object objobj and place objobj into/onto target locloc.
Task (Symbolic Form) OpenVLA Pi0 Pi0.5
Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env
Place(alphabet_soup,basket)Place(alphabet\_soup,basket) 1.00 0.97 0.00 0.96 0.00 0.00 0.98 0.98 0.00 0.88 0.00 0.06 0.96 0.96 0.00 0.94 0.00 0.68
Place(bbq_sauce,basket)Place(bbq\_sauce,basket) 0.98 0.90 0.00 0.94 0.00 0.00 0.98 1.00 0.00 1.00 0.00 0.46 1.00 0.98 1.00 1.00 0.02 0.96
$Place(butter,basket)$ 1.00 1.00 0.00 1.00 0.00 0.00 1.00 0.98 0.00 1.00 0.00 0.00 0.96 0.98 0.54 0.98 0.00 0.00
Place(chocolate_pudding,basket)Place(chocolate\_pudding,basket) 0.98 1.00 0.00 0.98 0.00 0.00 0.98 0.96 0.00 1.00 0.00 0.18 1.00 0.96 0.00 0.94 0.02 0.82
Place(cream_cheese,basket)Place(cream\_cheese,basket) 1.00 1.00 0.00 1.00 0.00 0.00 1.00 1.00 0.10 1.00 0.00 0.76 0.98 1.00 0.00 0.94 0.00 0.88
$Place(ketchup,basket)$ 1.00 1.00 0.00 1.00 0.00 0.00 0.96 0.84 0.00 0.92 0.00 0.24 0.98 0.96 0.20 0.96 0.02 0.92
$Place(milk,basket)$ 1.00 0.98 0.00 0.96 0.00 0.00 1.00 0.84 0.00 1.00 0.00 0.16 1.00 1.00 0.00 0.98 0.00 0.80
Place(orange_juice,basket)Place(orange\_juice,basket) 1.00 1.00 0.00 1.00 0.00 0.00 0.92 1.00 0.00 1.00 0.00 0.62 0.98 1.00 0.00 0.96 0.02 0.28
Place(salad_dressing,basket)Place(salad\_dressing,basket) 0.94 1.00 0.00 1.00 0.00 0.00 0.98 0.98 0.10 1.00 0.00 0.00 1.00 0.98 0.00 0.98 0.00 1.00
Place(tomato_sauce,basket)Place(tomato\_sauce,basket) 0.96 1.00 0.00 1.00 0.00 0.00 0.98 0.82 0.00 0.92 0.00 0.48 0.98 0.98 0.00 0.96 0.00 0.92
Average 0.99 0.98 0.00 0.98 0.00 0.00 0.98 0.94 0.00 0.90 0.00 0.29 0.98 0.98 0.17 0.96 0.01 0.73
5.3 Does the Model Generalize to New Objects?
(a) "salad dressing" in the training set.
(b) "salad dressing" not in the training set.
(c) Replace "salad dressing" with "alphabet soup".
Figure 4: Pick up salad dressing and place it in basket.
To investigate whether the VLA model exhibits true object-level generalization, we conducted tests by altering the manipulated object in the task “pick up the salad dressing and place it in the basket”. As shown in Figure 4, replacing the original salad dressing with an unseen but visually similar variant from outside the training set (second row) did not affect the model’s performance, suggesting a limited degree of generalization to perturbed object instances. However, when substituting salad dressing with an unrelated object (e.g., alphabet soup, third row), the model still executed the same action trajectory, attempting to pick up the new object despite its mismatch with the instruction. This behavior suggests that the model’s apparent generalization is largely driven by rote recall of action trajectories from training rather than genuine grounding of object identity. Moreover, its failure on semantically irrelevant objects highlights the inability to correctly align linguistic references with object semantics. Taken together, these observations indicate that the model’s performance reflects superficial pattern memorization rather than true object-level reasoning.
5.4 Does the Model Generalize to Varied Instructions?
(a) Original task.
(b) Semantic perturbation: Grab salad dressing and set it in basket.
(c) Meaningless task: fdsgfdsgsd or xxxxxx.
Figure 5: Pick up salad dressing and place it in basket.
We evaluated whether the VLA model robustly follows natural language instructions. As shown in Figure 5, the model successfully executed the original task (“pick up the salad dressing and place it in the basket”) and also completed the task when the instruction was reformulated with synonymous expressions (“grab the salad dressing and set it in the basket”). Although this outcome may appear to suggest robustness to linguistic variation, when the instruction was replaced with meaningless input (e.g., “xxx”), the model produced the same action sequence, again retrieving and executing the pattern of placing the salad dressing into the basket. This observation indicates that the apparent robustness does not arise from genuine semantic comprehension. Rather, the model relies predominantly on visual context and memorized action trajectories from training, while disregarding the linguistic content of the input. Consequently, the model’s consistent behavior across both paraphrased and nonsensical instructions should not be regarded as successful generalization to semantic perturbations, but instead as evidence of its inability to ground or interpret instructions, revealing an overdependence on rote pattern execution.
5.5 How Sensitive is the Model to Object Placement?
Figure 6: Success rates of OpenVLA, pi0, and pi0.5 under object position perturbations.
Although all three models achieve around 98% accuracy under the standard LIBERO, our perturbation experiments reveal significant robustness gaps that remain hidden in the original evaluation. As shown in Figure 6, the models are highly sensitive to the initial position of objects: for both OpenVLA and pi0, performance collapses once displacement exceeds 0.2 units, dropping to zero success, while pi0.5—benefiting from a stronger architecture and broader training data—exhibits improved resilience and sustains non-trivial accuracy up to 0.4 units before also collapsing to zero. These results indicate that high scores do not guarantee reliable performance in practice, as even small variations such as placing a cup in a slightly different position can cause complete task failure, underscoring the necessity of LIBERO-PRO to systematically reveal such robustness gaps and provide a more faithful comparison across models.
5.6 Is the Model Truly Understanding Tasks?
Figure 7: The model successfully executed two independent instructions, but failed to execute the combined instruction.
To assess whether the model demonstrates genuine task comprehension or merely memorizes action trajectories, we designed a sequence of three progressively structured tasks (as shown in Figure 7). In the first task, the agent is instructed to pick up the bowl and place it on the stove, and the model executes this sequence successfully, indicating its ability to reproduce simple manipulation behaviors. In the second task, the instruction is to turn on the stove, and again the model succeeds, suggesting competence in carrying out isolated atomic actions. However, in the third task, which composes these two operations into a longer sequence—first picking up the bowl, then placing it on the stove, and finally turning the stove on—the model consistently fails. Rather than demonstrating adaptive reasoning over the composed instruction, the model replays spurious trajectories resembling those seen during training, leading to unsuccessful execution. This discrepancy reveals a critical limitation: the model does not truly understand either the semantic goal of the task or the causal relationships between subtasks, but instead exhibits rote reproduction of memorized patterns from the training distribution, and the failure on even modestly composed instructions underscores the fragility of current approaches and highlights the need for evaluation protocols that probe beyond single-step or template-like behaviors.
6 Conclusion
In conclusion, our study reveals that the current LIBERO evaluation protocol is fundamentally flawed: by reusing identical training and evaluation tasks with only imperceptible perturbations, it fails to measure genuine capability. Reported accuracies above 90% largely reflect rote memorization of fixed mappings from the training set, rather than true comprehension of instructions or acquisition of robust action strategies. We call on the community to move beyond misleading evaluation practices and adopt our LIBERO-PRO benchmark, which provides a principled and effective framework for assessing the real generalization and understanding capabilities of VLA models.

### v2

@@
-LIBERO-PRO: Towards Robust and Fair Evaluation of Vision-Language-Action Models Beyond Memorization
+Xueyang Zhou1Yangming Xu1Guiyao Tie1Yongchao Chen2Guowen Zhang1 Duanfeng Chu3 Pan Zhou1Lichao Sun4
 Abstract
@@
 Our code is available at: https://github.com/Zxy-MLlab/LIBERO-PRO.
-Figure 1: Model trajectory consistency under object and instruction modifications.
-Original task” denotes the standard evaluation, while “Replace,” “Change Pos.,” “Remove,” and “Messy instruction” introduce perturbations to the target object or instruction.
-Across all settings, the model yields nearly identical trajectories, suggesting a lack of genuine task understanding and environmental perception.
+Keywords Vision-Language-Action $\cdot$ Embodied AI $\cdot$ Benchmark $\cdot$ Robustness Evaluation
+Figure 1: VLA models evaluated on LIBERO can be broadly grouped into five families according to their action decoding strategies: autoregressive, continuous, flow/diffusion, hybrid, and adapter-based methods.
+Although these different paradigms all achieve over 70% performance on LIBERO, their performance drops substantially on the more challenging LIBERO-PRO benchmark.
+This suggests that, despite strong results on LIBERO, current VLA models still suffer from limited robustness and generalization under more demanding evaluation settings.
 1 INTRODUCTION
-Vision-Language-Action (VLA) models have recently emerged as a central paradigm for bridging perception, language understanding and action execution in embodied agentsSapkota et al.
-(2025).
-A crucial factor in advancing this line of research is the availability of standardized benchmarksJames et al.
-(2020); Liu et al.
-(2023); Mees et al.
-(2022); Li et al.
-(2024b); Nasiriany et al.
-(2024); Walke et al.
-(2023) that allow fair and reproducible comparison across different methods.
-Among them, LIBEROLiu et al.
-(2023) has rapidly become the most widely adopted evaluation suite for VLA, serving as the de facto standard for reporting performance.
-As shown in Table 1, models ranging from large-scale pre-trained to specialized systems are all benchmarked on LIBERO, making it the “common currency” of evaluation in this field.
-Consequently, the reliability of LIBERO as an evaluation protocol plays a decisive role in shaping research directions in VLA.
-Figure 2: Success rates of OpenVLA, pi0, pi0.5, and univla under object position perturbations.
-However, the current LIBERO training–evaluation setup may suffer from serious limitations.
-In fact, the evaluation tasks are exactly the same as the training tasks, with the only difference being slight perturbations in the initial states of objects—perturbations that are almost invisible to the human eye.
-As a result, if a model overfits the training set, it can easily achieve high scores during evaluation, but its accuracy drops to nearly zero under minor changes in object initial positions or task instructions, as shown in Figure 2.
-Yet such results do not indicate that the model has learned to understand the environment or the task.
-On the contrary, we find that models rely on fixed mappings from the training tasks.
-As illustrated in Figure 1, when we introduced four types of modifications—(1) replacing the target object with an unrelated one; (2) relocating its initial position; (3) removing it entirely; and (4) corrupting the instruction with nonsensical tokens—their trajectories remained nearly identical.
-This outcome shows that, despite achieving over 90% on the standard LIBERO benchmark, the reported scores reflect rote recall rather than genuine task understanding.
-Obviously, the current evaluation protocol is inadequate and fails to effectively assess model architectures and algorithms.
-To address the above shortcomings, we systematically propose an extended evaluation suite, i.e., LIBERO-PRO, to enable a more comprehensive and reasonable assessment of VLA models.
-We design evaluation scenarios along four key dimensions: manipulated objects, initial states, task instructions, and environments.
-While ensuring that each task remains coherent and executable, we introduce carefully designed perturbations to the original LIBERO tasks to simulate real-world situations.
-Using this suite, we evaluate state-of-the-art models including OpenVLAKim et al.
-(2024) and pi0/0.5, and the results show that they exhibit near-complete failure under several of these settings: the models cannot cope with changes in object positions, nor can they correctly follow slightly rephrased task instructions—even when such instructions have appeared in other training contexts.
-These findings reveal that the commonly reported accuracies of over 90% are largely artifacts of task memorization, rather than evidence of genuine perception or task understanding.
-We therefore call on the community to adopt evaluation standards that incorporate such necessary perturbations, ensuring that reported progress reflects capabilities required for practical use.
-Our main contributions are as follows:
-  • Through extensive experiments, we demonstrate that the current LIBERO evaluation protocol is fundamentally flawed: high scores primarily reflect rote memorization of training data rather than genuine task understanding or execution ability.
-  • We introduce LIBERO-PRO, a plug-and-play benchmark that incorporates carefully designed perturbations across four dimensions—manipulated objects, initial states, task instructions, and environments—and supports randomized combinations of these dimensions for more faithful evaluation.
-  • We benchmark state-of-the-art models, including OpenVLA, pi0 and pi0.5, on LIBERO-PRO and show that they perform poorly, highlighting the inadequacy of the current LIBERO protocol and calling for LIBERO-PRO as a fairer and more reliable evaluation standard.
-2 Related Work
-Vision-Language-Action Models.
-Vision-Language-Action (VLA) models have rapidly emerged as a central paradigm for embodied AI, aiming to unify perception, natural language understanding, and action execution within a single framework.
-Recent years have witnessed an explosion of approaches, ranging from large-scale pre-trained models that leverage multimodal corpora to task-specific architectures tailored for robotic manipulation and navigationMa et al.
-(2025); Li et al.
-(2025b); Din et al.
-(2025).
-While these efforts have significantly advanced the field, they also underscore the urgent need for standardized evaluation benchmarksLiu et al.
-(2023); Mees et al.
-(2022); James et al.
-(2020).
-Without such benchmarks, it becomes difficult to conduct fair and reproducible comparisons across methods, hindering the cumulative progress of VLA research.
-Benchmarks for VLA Evaluation.
-A number of benchmarks have been developed to enable systematic evaluation of VLA models, including RLBench James et al.
-(2020), CALVIN Mees et al.
-(2022), RoboCasa Nasiriany et al.
-(2024), and BridgeData Walke et al.
-(2023).
-These benchmarks provide valuable testbeds for task execution under vision and language guidance, yet their adoption has remained relatively limited within the broader community.
-In contrast, LIBERO Liu et al.
-(2023) has quickly emerged as the dominant benchmark and is now the most widely adopted protocol for evaluating VLA models.
-Its standardized task suite and unified reporting metrics have made it the de facto standard for performance comparison, to the extent that virtually all recent VLA studies report results on LIBERO.
-As such, LIBERO not only serves as the “common currency” of evaluation in this field, but also exerts a strong influence on research directions and claims of progress.
-Table 1: Incomplete statistics of VLA models evaluated on LIBERO.
-Numbers shown to two decimal places.
-‘–’ denotes not reported.
-Single: One checkpoint trained for all tasks.
-Multi: Separate checkpoint trained for each task.
-Model Year Size Method Single/Multi Goal Object Spatial 10 Average
-pi0.5 Intelligence et al. (2025) 2025 $\sim$7B SFT Single 0.98a 0.98 0.98a 0.92 0.97
-CogACT Li et al. (2024a) 2024 $\sim$7B SFT - 0.90 0.98b 0.97b 0.89 0.98a
-GeoVLA Sun et al. (2025) 2025 $\sim$7B SFT - 0.97b 0.99a 0.98a 0.97a 0.98a
-OpenVLA Kim et al. (2024) 2024 $\sim$7B SFT Multi 0.98a 0.99a 0.98a 0.93b 0.97b
-pi0 Black et al. (2024) 2024 $\sim$7B SFT Single 0.92 0.98b 0.97b 0.96b
-UniVLA* Wang et al. (2025b) 2025 $\sim$8B SFT Single 0.94 0.99a 0.95 0.94b 0.96b
-SwitchVLA Li et al. (2025c) 2025 SFT Single 0.93 0.93
-TGRPO Chen et al. (2025) 2025 $\sim$7B RL Multi 0.91 0.91
-GraspVLA Deng et al. (2025) 2025 $\sim$2B SFT Single 0.91 0.94 0.82 0.89
-UniVLA Bu et al. (2025) 2025 $\sim$7B SFT Multi 0.90 0.94 0.91 0.79 0.89
-TriVLA Liu et al. (2025) 2025 $\sim$3B SFT - 0.90 0.94 0.91 0.73 0.87
-MolmoAct Lee et al. (2025) 2025 $\sim$7B SFT Multi 0.88 0.95 0.87 0.77 0.87
-CronusVLA Li et al. (2025a) 2025 $\sim$7B SFT Multi 0.91 0.95 0.90 0.69 0.86
-SmolVLA Shukor et al. (2025) 2025 $\sim$2B SFT Single 0.91 0.94 0.93 0.73 0.86
-ThinkAct Huang et al. (2025) 2025 $\sim$7B SFT+RL - 0.87 0.91 0.88 0.71 0.84
-World-VLA Cen et al. (2025) 2025 SFT Multi 0.83 0.96 0.88 0.60 0.82
-Cot-VLA Zhao et al. (2025) 2025 $\sim$7B SFT - 0.88 0.92 0.88 0.69 0.81
-VLA-RL Lu et al. (2025) 2025 $\sim$7B SFT+RL Multi 0.82 0.92 0.90 0.60 0.81
-SpatialVLA Qu et al. (2025) 2025 $\sim$4B SFT Multi 0.79 0.90 0.88 0.56 0.78
-Octo Team et al. (2024) 2024 $\sim$93M SFT Multi 0.85 0.86 0.79 0.51 0.75
-VQ-VLA Wang et al. (2025c) 2025 $\sim$7B SFT Single 0.75 0.60 0.68
-  • aBest (cell shaded light blue).   bSecond-best (cell shaded light green).
-Limitations of Current Evaluation Practices.
-Several recent studies have noted that the high scores reported on LIBERO often fail to translate into reliable performance, even on tasks with only minor variations.
-This has raised serious concerns about the validity of its evaluation protocolZhou et al.
-(2025); Fang et al.
-(2025); Li (2025).
-In response, some researchers have explored stronger algorithms or proposed stress tests Gu et al.
-(2025); Guruprasad et al.
-(2024); Valle et al.
-(2025); Kube et al.
-(2025); Wang et al.
-(2025a)to encourage more robust VLA systems.
-However, the field still lacks a systematic and standardized framework for robust evaluation under LIBERO.
-The majority of works continue to rely on the original, flawed setup, reporting numbers that may not reflect genuine capability.
-To address this gap, we propose LIBERO-PRO, a plug-and-play extension of LIBERO designed to foster fair comparison across models and to advance research on building VLA systems that can withstand reasonable and practically necessary perturbations.
+Vision-Language-Action (VLA) models have recently emerged as a fundamental paradigm for bridging perception, language understanding, and action execution in embodied agents [17].
+A central catalyst for progress in this domain is the development of standardized benchmarks [7, 13, 15, 12, 16, 19], which enable reproducible, fair, and quantitative comparison across competing methods.
+Among these, LIBERO [13] has rapidly become the most widely adopted evaluation suite for VLAs, functioning as the de facto standard for performance reporting.
+As summarized in Figure 6, both large-scale pre-trained models and task-specific systems are benchmarked on LIBERO, rendering it the “common currency” for evaluation in this field.
+Consequently, the methodological rigor and reliability of LIBERO directly influence how research directions in VLA are defined and pursued.
+Despite its influence, the current LIBERO training–evaluation configuration exhibits fundamental limitations.
+Specifically, the evaluation tasks are identical to the training tasks, differing only by marginal perturbations in initial object states—variations so subtle as to be visually imperceptible.
+As a result, models that overfit to the training distribution can still achieve near-perfect evaluation scores, while their accuracy collapses under even minor changes in object position or task phrasing (Figure 6).
+Such inflated performance fails to reflect genuine task comprehension or environmental reasoning.
+Our analysis further reveals that VLA models often rely on memorized task mappings rather than learned representations.
+Specifically, when we applied four systematic modifications—(1) replacing the target object with an unrelated one, (2) relocating its initial position, (3) removing it entirely, and (4) corrupting the instruction with nonsensical tokens—the resulting trajectories remained nearly unchanged.
+These findings demonstrate that, despite achieving over 90% success on the standard LIBERO benchmark, current results primarily indicate rote recall rather than robust task understanding.
+Hence, the existing evaluation protocol is insufficient to rigorously assess the reasoning, compositionality, and generalization capacity of modern VLA architectures.
+To address these shortcomings, we introduce an extended evaluation suite, LIBERO-PRO, designed to provide a more comprehensive and reliable measure of VLA performance.
+LIBERO-PRO systematically constructs evaluation scenarios along four orthogonal dimensions—manipulated objects, initial states, task instructions, and environments—while ensuring that each perturbed task remains logically coherent and physically executable.
+Through carefully controlled perturbations, we simulate realistic variations that better approximate the diversity of real-world conditions.
+Using this suite, we evaluate state-of-the-art VLA models, including OpenVLA [8] and Pi0/Pi0.5, and observe near-complete performance collapse under several perturbation regimes: models fail to adapt to altered object positions and exhibit severe sensitivity to even minimal instruction paraphrasing, despite encountering semantically equivalent expressions during training.
+These findings reveal that the widely reported accuracies exceeding 90% largely reflect memorization rather than genuine perceptual grounding or task-level reasoning.
+We therefore advocate for evaluation protocols that incorporate structured perturbations, ensuring that reported progress aligns with the requirements of robust and generalizable embodied intelligence.
+Our key contributions are summarized as follows.
+  • We empirically demonstrate that the current LIBERO evaluation protocol substantially overestimates model competence, as its high reported scores largely stem from memorization of training configurations rather than genuine understanding or transferable reasoning ability.
+  • We introduce LIBERO-PRO, a plug-and-play benchmark that integrates systematically designed perturbations across four core dimensions—manipulated objects, initial states, task instructions, and environments—supporting both single- and multi-dimensional randomized combinations for rigorous evaluation.
+  • We benchmark leading VLA models, including OpenVLA, Pi0, and Pi0.5, on LIBERO-PRO and reveal significant performance degradation, underscoring the limitations of the original LIBERO protocol and motivating LIBERO-PRO as a fairer, more faithful, and reproducible evaluation standard.
+2 Observing Overfitting Issues
+Despite most VLA models achieving over 95% success on LIBERO, our qualitative experiments reveal several systematic failure patterns that collectively indicate severe overfitting to the training distribution.
+Specifically, we observe: (1) brittle visual perception, manifesting as object or location hallucinations; (2) insensitivity to instruction semantics, where actions are driven by trajectory recall rather than language grounding; and (3) a lack of compositionality, with models unable to integrate individually learned subtasks into coherent multi-step behaviors.
+2.1 Visual perceptual hallucinations
+(a) "salad dressing" in the training set.
+(b) "salad dressing" not in the training set.
+(c) Replace "salad dressing" with "alphabet soup".
+Figure 2: Pick up salad dressing and place it in basket.
+Executing the grasp on an unrelated object.
+As shown in Figure 2(b), when we replace the target salad dressing bottle with an unrelated item—alphabet soup—the model still executes the instruction “Pick up the salad dressing and place it in the basket” by reaching toward and grasping the alphabet soup.
+This behavior indicates that the model does not reliably bind the instruction to the correct visual entity.
+Instead, it appears to infer the intended action from the overall scene context and subsequently replay a familiar trajectory, regardless of whether the visual evidence supports the action.
+Executing the grasp despite the object’s removal.
+A similar error emerges when the target object is entirely removed from the scene.
+As illustrated in Figure 2(c), even though the salad dressing is absent, the model still performs the grasping motion toward the location where the object typically appears.
+This “grasping the air” behavior further indicates that the policy does not perform online visual verification of object existence before initiating the action.
+2.2 Instruction understand hallucinations
+Ignoring meaningless instructions.
+(a) Original task.
+(b) Meaningless task: fdsgfdsgsd or xxxxxx.
+Figure 3: Pick up salad dressing and place it in basket.
+Figure 3(b), when we replace the original instruction with a meaningless character sequence (e.g., “fdsgfdsgsd” or “xxx”), the model still executes nearly the same action trajectory as in the original task “Pick up the salad dressing and place it in the basket.” Rather than recognizing the absence of valid semantic content, the model proceeds to grasp the salad dressing and attempt the placement action without any hesitation.
+This behavior indicates that the policy is largely insensitive to linguistic semantics and that the instruction input contributes little to the decision-making process.
+Instead, the model appears to rely primarily on visual context to retrieve a memorized visuomotor pattern, further reinforcing that its apparent instruction-following capability does not stem from genuine language grounding.
+2.3 Location perception hallucinations
+Figure 4: Success rates of OpenVLA, pi0, and pi0.5 under object position perturbations.
+Our qualitative results show that the policy is highly sensitive to the object’s initial position: when the manipulated object is displaced away from the coordinates commonly observed in the training demonstrations, the model consistently fails to execute the grasp.
+Even small offsets lead the end-effector to move toward the original, memorized location rather than the object’s actual position.
+This behavior indicates that the policy does not perform online spatial reasoning or object localization.
+As shown in Figure 4, the models are extremely sensitive to shifts in the object’s initial position: for both OpenVLA and pi0, task success collapses once the displacement exceeds 0.2 units, dropping sharply to zero thereafter.
+Even the stronger pi0.5 model—benefiting from a larger architecture and more diverse training data—only maintains non-trivial performance up to roughly 0.4 units before also degrading to zero.
+2.4 Lack of task understanding
+Figure 5: The model successfully executed two independent instructions, but failed to execute the combined instruction.
+To evaluate whether the model possesses genuine task-level understanding rather than merely reproducing isolated demonstrations, we design a sequence of three progressively structured tasks (Figure 5).
+In the first task—pick up the bowl and place it on the stove—the model succeeds, indicating it can replicate a simple manipulation trajectory.
+In the second task—turn on the stove—the model again performs reliably, suggesting competence with single-step atomic actions.
+However, when the two subtasks are combined into a single multi-step instruction—pick up the bowl, place it on the stove, and then turn on the stove—the model consistently fails.
+Rather than assembling the subtasks into a coherent execution plan, it replays partial or spurious trajectories reminiscent of the training distribution.
+This discrepancy reveals that the model lacks compositional reasoning: it does not understand the temporal or causal structure linking the subtasks, but instead depends on memorized action fragments.
+Even such modest task compositions exceed its generalization capacity, highlighting the fragility of current VLA models.
+These observations suggest that the strong in-distribution performance of current Visual-Language-Action (VLA) models may rely heavily on superficial regularities in the training data.
+Although a VLA policy is intended to learn a mapping $\pi:(\mathcal{O},\mathcal{L})\rightarrow\mathcal{A}$ from visual observations and language instructions to actions, our results indicate that the learned policy often fails to robustly ground actions in the actual observation-instruction pair.
+Instead, it appears to exploit shortcuts such as canonical scene layouts, recurring object appearances, typical object positions, and repeated motion trajectories.
+This provides a plausible explanation for the observed failures: the model grasps absent or incorrect objects, ignores meaningless instructions, reaches toward memorized locations under positional perturbations, and fails to compose individually learned subtasks into coherent multi-step behaviors.
+Overall, these results suggest that current VLA models may still depend more on shortcut feature-action correlations than on robust and compositional visuomotor understanding.
 3 Problem Definition
 This section formalizes the Vision-Language-Action (VLA) task framework and examines limitations in the LIBERO benchmark’s evaluation protocol that may obscure true task competence.
-3.1 Formal Task Definition
-We define a VLA task to establish a rigorous basis for evaluating robotic systems.
-$\tau=(\ell,\mathcal{O},\mathcal{E},p_{0},G),$ (1)
-where $\ell$ is a natural language instruction specifying the task objective, $\mathcal{O}$ is a finite set of objects with attributes (e.g., shape, color, size, initial pose), $\mathcal{E}$ denotes the static environmental context (e.g., scene geometry, furniture layout), $p_{0}$ is a probability distribution over initial states $s_{0}$ (encoding init position and robot proprioception), and $G(s)$ is a binary goal predicate, returning $1$ if state $s$ satisfies $\ell$, and 0 otherwise.
-The performance of a policy $\pi_{\theta}(a\mid s,\ell)$, parameterized by $\theta$, is measured by the expected success probability over a finite horizon $T$:
-$\mathbb{P}_{s_{0}\sim p_{0},\pi_{\theta}}\left[G(s_{T})=1\right],$ (2)
-where $s_{T}$ is the terminal state after $T$ steps. This formulation captures the integration of perception, language understanding, and action planning in VLA systems.
+3.1 Formal VLA Task Definition
+We formalize a vision-language-action (VLA) task instance as
+$\mathcal{T}=(l,e,G),$ (1)
+where $l\in\mathcal{L}$ is a natural-language instruction, $e\in\mathcal{E}$ denotes an environment instance, $\mathcal{E}$ is the space of environment instances, and $G:\mathcal{S}\times\mathcal{L}\to\{0,1\}$ is a binary success predicate indicating whether a state satisfies the task specified by the instruction.
+An environment instance is defined as
+$e=(\mathcal{W},\mathcal{O},R_{0}),$ (2)
+where $\mathcal{W}$ denotes the visual environment context (e.g., background, textures, and lighting conditions), $\mathcal{O}$ denotes the set of interactive objects in the scene, and $R_{0}$ denotes the initial spatial configuration of the objects.
+Let $s_{0}$ denote an initial state consistent with $R_{0}$.
+A policy $\pi_{\theta}(a_{t}\mid o_{t},l)$ induces trajectories according to
+$a_{t}\sim\pi_{\theta}(\cdot\mid o_{t},l),\qquad s_{t+1}\sim P(\cdot\mid s_{t},a_{t}),$ (3)
+where $o_{t}$ denotes the agent observation at time $t$, and $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ denotes the environment dynamics.
+We evaluate a policy by its success probability within a finite horizon $H$:
+$J(\pi_{\theta};\mathcal{T})=\mathbb{P}_{\pi_{\theta},\,e}\!\left[\exists t\in\{0,\dots,H\},\;G(s_{t},l)=1\right],$ (4)
+where the probability is taken over trajectories induced by $\pi_{\theta}$ under the environment $e$. This objective captures the joint requirements of language grounding, perception, and action.
 3.2 Current LIBERO Evaluation Protocol
-The LIBERO benchmark, while valuable for VLA research, employs an evaluation protocol that may overestimate model competence due to its homogeneous train/test split.
-Each test task $\tau^{\prime}\in\mathcal{T}_{\mathrm{test}}$ mirrors a training task $\tau\in\mathcal{T}_{\mathrm{train}}$:
-$\ell^{\prime}=\ell,\quad\mathcal{O}^{\prime}=\mathcal{O},\quad\mathcal{E}^{\prime}=\mathcal{E},\quad G^{\prime}=G,$ (3)
-with the only difference being a minor perturbation in the initial state distribution:
-$d_{\mathrm{TV}}(p_{0},p_{0}^{\prime})\leq\epsilon,\quad\epsilon\ll 1.$ (4)
-Due to the small $\epsilon$, test tasks closely resemble training tasks, allowing policies to achieve high performance through memorization rather than robust generalization.
-This homogeneity may inflate performance metrics, masking limitations in compositional generalization and adaptability to diverse real-world scenarios, a critical concern for embodied AI.
-Figure 3: LIBERO-PRO benchmark task overview.
-We extend the four original task categories in LIBERO by introducing object attribute perturbations, initial position perturbations, instruction perturbations, and environmental perturbations.
+While LIBERO has been an influential benchmark for VLA research, its current evaluation protocol exhibits limited distributional shift between training and test tasks.
+Under our formulation, each task is specified by an instruction $l$, an environment instance $e=(\mathcal{W},\mathcal{O},R_{0})$, and a success predicate $G$.
+In LIBERO, a test task $\mathcal{T}^{\prime}\in\mathcal{T}_{\mathrm{test}}$ is typically constructed from a corresponding training task $\mathcal{T}\in\mathcal{T}_{\mathrm{train}}$ such that
+$l^{\prime}=l,\qquad\mathcal{W}^{\prime}=\mathcal{W},\qquad\mathcal{O}^{\prime}=\mathcal{O},\qquad G^{\prime}=G,$ (5)
+with variation arising only from small perturbations to the initial spatial configuration $R_{0}$, i.e., $R_{0}^{\prime}$ differs from $R_{0}$ only by minor perturbations.
+Consequently, test tasks remain close to training tasks, and evaluation primarily reflects robustness to small variations in object configurations rather than generalization to novel task compositions or environments.
+As a result, strong performance may largely arise from interpolation within a narrow neighborhood of the training distribution, potentially overestimating a model’s ability to generalize under broader distribution shifts.
 4 Methodology
 4.1 Proposed Perturbation-Based Evaluation Framework
-For a more reasonable evaluation of VLA models, we propose a framework that introduces controlled perturbations across four orthogonal dimensions, manipulated objects, initial states, task instructions, and environments.
-For a task $\tau=(\ell,\mathcal{O},\mathcal{E},p_{0},G)$, we generate perturbed tasks $\tau^{(k)}=\phi_{k}(\tau)$ using operators $\phi_{k}$, $k\in\{\mathrm{O},\mathrm{S},\mathrm{L},\mathrm{E}\}$, defined as follows:
-Object Attribute Perturbation ($k=\mathrm{O}$): Modifies non-essential object attributes, such as color, texture, or size, while preserving semantic equivalence, e.g., changing a red cup to a blue cup.
-This tests the model’s robustness to superficial visual changes critical for real-world adaptability.
-The perturbed task is:
-$\tau^{(\mathrm{O})}=\left(\ell,\mathcal{O}^{(\mathrm{O})},\mathcal{E},p_{0},G\right).$ (5)
-Initial Position Perturbation ($k=\mathrm{I}$): Alters initial object positions to modify absolute and relative arrangements, ensuring physical plausibility, e.g., moving a cup relative to a plate.
-This probes the model’s spatial reasoning under varied layouts.
-The perturbed task is:
-$\tau^{(\mathrm{I})}=\left(\ell,\mathcal{O},\mathcal{E},p_{0}^{(\mathrm{I})},G\right).$ (6)
-Instruction Perturbation ($k=\mathrm{L}$): Perturbations are applied at both the semantic and task levels.
-Semantic perturbations adjust the linguistic expression while preserving the original task intent, such as paraphrasing “pick up the bottle” to “grab the bottle” or substituting with a jar.
-This evaluates the agent’s understanding of language and semantics.
-The perturbed task is:
-$\tau^{(\mathrm{L})}=\left(\ell^{(\mathrm{L})},\mathcal{O},\mathcal{E},p_{0},G\right),$ (7)
-Task perturbations reasonably modify the task itself by replacing the target object or altering the action, e.g., changing “pick up the bottle” to “pick up the bowl.” We ensure that all actions and objects appear in the training set to support fair and meaningful generalization.
-The perturbed task is:
-$\tau^{(\mathrm{L})}=\left(\ell^{(\mathrm{L})},\mathcal{O}^{(\mathrm{L})},\mathcal{E},p_{0},G^{(\mathrm{L})}\right).$ (8)
-Environment Perturbation ($k=\mathrm{E}$): Modifies the background’s appearance, e.g., changing a workspace, without affecting task feasibility, to assess robustness to environmental variations.
-The perturbed task is:
-$\tau^{(\mathrm{E})}=\left(\ell,\mathcal{O},\mathcal{E}^{(\mathrm{E})},p,G\right).$ (9)
+Figure 6: Overview of LIBERO-PRO.
+LIBERO-PRO extends LIBERO with five perturbation types: object, configuration, instruction, task, and environment.
+These perturbations vary object attributes, object placements, language descriptions, task goals/object sets, and visual context, respectively, enabling systematic evaluation under diverse distribution shifts.
+To ensure feasibility, task perturbations are excluded from cross-type combinations, while the other perturbation types can be freely composed.
+To enable more principled evaluation of VLA models, we design a perturbation-based framework that introduces controlled variations across four orthogonal dimensions: object attributes, initial spatial configurations, task instructions, and environments.
+Given a task $\mathcal{T}=(l,e,G)$ with environment instance $e=(\mathcal{W},\mathcal{O},R_{0})$, we generate perturbed variants $\mathcal{T}^{(k)}=\phi_{k}(\mathcal{T})$, where $k\in\{\mathrm{O},\mathrm{R},\mathrm{L},\mathrm{E}\}$.
+Object attribute perturbation ($k=\mathrm{O}$) modifies non-essential visual attributes of objects, such as color, texture, or size, while preserving semantic equivalence. The resulting task is
+$\mathcal{T}^{(\mathrm{O})}=(l,e^{(\mathrm{O})},G),\qquad e^{(\mathrm{O})}=(\mathcal{W},\mathcal{O}^{(\mathrm{O})},R_{0}).$
+Initial configuration perturbation ($k=\mathrm{R}$) alters the absolute and relative positions of objects while maintaining physical plausibility. The perturbed task is
+$\mathcal{T}^{(\mathrm{R})}=(l,e^{(\mathrm{R})},G),\qquad e^{(\mathrm{R})}=(\mathcal{W},\mathcal{O},R_{0}^{(\mathrm{R})}).$
+Instruction perturbation ($k=\mathrm{L}$) includes both semantic- and task-level variations. Semantic perturbations rephrase the instruction while preserving task intent:
+$\mathcal{T}^{(\mathrm{L}_{\mathrm{sem}})}=(l^{(\mathrm{L})},e,G).$
+Task perturbation ($k=\mathrm{T}$) modifies the target object or action while remaining within the training distribution. The perturbed task is
+$\mathcal{T}^{(\mathrm{T})}=(l,e,G^{(\mathrm{T})}).$
+Environment perturbation ($k=\mathrm{E}$) modifies the visual environment context, such as background, lighting, or texture, without altering task feasibility. The resulting task is
+$\mathcal{T}^{(\mathrm{E})}=(l,e^{(\mathrm{E})},G),\qquad e^{(\mathrm{E})}=(\mathcal{W}^{(\mathrm{E})},\mathcal{O},R_{0}).$
 4.2 Perturbation Constraints
-To ensure each perturbed task $\tau^{(k)}$ remains semantically valid, we impose two constraints. First, perturbations are confined within a dimension-specific neighborhood:
-$x^{(k)}\in\mathcal{N}_{k}(x),\quad\mathrm{dist}_{k}(x,x^{(k)})\leq\delta_{k},$ (10)
-where $x$ is the original component, $x^{(k)}$ is its perturbed version, and $\delta_{k}$ controls the magnitude of perturbation for dimension $k\in\{\mathrm{O},\mathrm{S},\mathrm{L},\mathrm{E}\}$.
-Second, we enforce non-trivial task-level separation:
-$d_{\mathrm{TV}}(\tau,\tau^{(k)})>\epsilon,\quad\forall k\in\{\mathrm{O},\mathrm{S},\mathrm{L},\mathrm{E}\},$ (11)
-where $d_{\mathrm{TV}}$ is the total variation distance, ensuring that perturbations lead to substantive deviations.
-Multiple perturbed tasks are generated per category by applying $\phi_{k}$, with $\delta_{k}$ varied to characterize the sensitivity of policies under different perturbation magnitudes.
+To ensure that each perturbed task remains semantically coherent and executable, we impose two classes of constraints. First, perturbations are bounded within a dimension-specific neighborhood:
+$x^{(k)}\in\mathcal{N}_{k}(x),\qquad\mathrm{dist}_{k}(x,x^{(k)})\leq\delta_{k},$
+where $x$ denotes a task component, $x^{(k)}$ its perturbed counterpart, and $\delta_{k}$ controls the perturbation magnitude along dimension $k$. Second, we enforce non-trivial task-level separation:
+$d_{\mathrm{TV}}(\mathcal{T},\mathcal{T}^{(k)})>\epsilon,$
+where $d_{\mathrm{TV}}$ denotes the total variation distance, ensuring that perturbations induce meaningful behavioral changes.
+By varying $\delta_{k}$, we generate multiple perturbed variants for each category and characterize policy sensitivity to perturbation intensity.
 4.3 LIBERO-PRO Evaluation Benchmark
-Building on these principles, we introduce LIBERO-PRO, a plug-and-play evaluation benchmark that enables systematic assessment under multiple dimensions of variation—object, position, instruction, task, and environment—as well as their arbitrary combinations, as shown in Figure 3.
-Specifically, for object variation, we create new assets for the four LIBERO tasks by modifying the appearance, size, and color of the original objects.
-For position variation, we define alternative spatial regions for all manipulable objects while strictly adhering to physical constraints and task definitions.
-For instruction variation, we provide three distinct paraphrased versions of each task description.
-For task variation, we redesign tasks in a principled manner to ensure feasibility, and supply new sets of relevant objects and target states for each modified task.
-For environment variation, we leverage the five environment categories built into the LIBERO benchmark and randomly substitute them across tasks.
-To avoid meaningless combinations, we restrict task variation from being combined with other types, while allowing all other dimensions of variation to be freely and randomly combined for more comprehensive evaluation.
+Building upon the above framework, we introduce LIBERO-PRO, a plug-and-play benchmark for systematic evaluation across multiple dimensions of variation, including object attributes, spatial configurations, instructions, tasks, and environments, as well as their combinations.
+For object variation, we modify object color, texture, and scale.
+For configuration variation, we alter object placements while preserving physical validity.
+For instruction variation, we generate paraphrased task descriptions.
+For task variation, we redesign tasks to introduce new object sets and goals while maintaining executability.
+For environment variation, we vary visual context such as backgrounds and lighting.
+To avoid infeasible combinations, task-level perturbations are excluded from cross-type compositions, while the remaining perturbation types may be freely combined to construct diverse evaluation settings.
 5 Experiments
 5.1 Settings
+Figure 7: Main results on LIBERO-PRO.
+Despite high average success on standard LIBERO tasks, current VLA models fail sharply under controlled perturbations, particularly position and task shifts.
+Robustness to object, semantic instruction, and environment changes is limited and varies across models and task suites.
 We conduct experiments on LIBERO-PRO, which applies controlled perturbations from four perspectives: object properties, initial positions, task instructions (covering both semantic and structural variations), and environments.
-We evaluate three representative models—OpenVLAKim et al.
-(2024), pi0Black et al.
-(2024), and pi0.5Intelligence et al.
-(2025).
+We evaluate three representative models—OpenVLA[8], pi0[1], and pi0.5[6].
 For fairness, OpenVLA is assessed using four official checkpoints trained on different LIBERO suites, whereas pi0 and pi0.5 are evaluated with a single official checkpoint.
@@
 5.2 Main Results
-We conducted a comprehensive evaluation of state-of-the-art vision-language-action (VLA) models Black et al.
-(2024); Kim et al.
-(2024) using the LIBERO-PRO evaluation suite, which introduces perturbations across four dimensions: objects, goals, spatial configurations, and task sets (see Tables 2, 3, 4, 5).
+We conducted a comprehensive evaluation of state-of-the-art vision-language-action (VLA) models [1, 8] using the LIBERO-PRO evaluation suite, which introduces perturbations across four dimensions: objects, goals, spatial configurations, and task sets (see Tables 7).
 Our findings are as follows: (1) Failure under simple perturbations.
@@
 Overall, these results show that current VLA models, despite excelling on the standard LIBERO, lack true robustness and generalization—highlighting the need for rigorous evaluation frameworks.
-Table 2: Task-wise LIBERO and LIBERO-PRO performance of three VLA models on the libero-goal benchmark.
-Accuracy on test (Ori: origina, Obj: object, Spa: spatial, Sem: semantic, Task: task-level, Env: environment).
-Action notation: $Open(x,y)$ = open target $y$ of container $x$; $Put(obj,loc)$ = place object objobj onto/into location locloc; $Push(obj,loc)$ = push object objobj toward location locloc; $TurnOn(obj)$ = activate object objobj.
-Task (Symbolic Form) OpenVLA Pi0 Pi0.5
-Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env
-$Open(cabinet,drawer_{mid})$ 0.98 0.96 0.00 1.00 0.00 0.96 0.92 0.94 0.00 0.94 0.00 0.00 0.96 0.96 0.00 0.94 0.04 0.92
-$Put(bowl,drawer_{top})$ 0.88 0.88 0.00 0.96 0.00 0.82 0.76 0.86 0.00 0.94 0.00 0.10 0.98 0.98 0.94 1.00 0.02 1.00
-$Push(plate,stove_{front})$ 1.00 1.00 0.00 1.00 0.00 1.00 0.94 0.93 0.00 0.94 0.00 0.54 0.96 0.98 0.00 0.98 0.00 0.28
-$Put(bowl,plate)$ 1.00 1.00 0.00 1.00 0.00 1.00 1.00 0.99 0.00 0.96 0.00 0.96 0.90 0.90 0.00 0.92 0.02 0.04
-$Put(bowl,stove)$ 0.94 1.00 0.00 1.00 0.00 1.00 0.98 0.97 0.00 0.98 0.00 0.98 0.98 0.96 0.00 0.98 0.04 0.78
-$Put(bowl,cabinet_{top})$ 1.00 0.98 0.00 1.00 0.00 1.00 0.98 0.94 0.00 0.92 0.00 0.94 0.96 0.96 0.00 0.96 0.02 0.00
-Put(cream_cheese,bowl)Put(cream\_cheese,bowl) 1.00 1.00 0.00 1.00 0.00 1.00 0.96 0.96 0.00 0.96 0.00 0.46 0.98 1.00 0.98 0.98 0.02 0.94
-Put(wine_bottle,rack)Put(wine\_bottle,rack) 1.00 1.00 0.00 1.00 0.00 1.00 0.68 0.84 0.00 0.64 0.00 0.00 1.00 0.98 0.88 0.96 0.02 0.12
-$Put(wine\_bottle,cabinet_{top})$ 0.96 0.83 0.00 0.88 0.00 0.88 0.94 1.00 0.00 1.00 0.00 0.00 1.00 1.00 0.98 0.98 0.02 0.48
-$TurnOn(stove)$ 1.00 1.00 0.00 1.00 0.00 1.00 1.00 1.00 0.00 1.00 0.00 0.00 1.00 1.00 0.00 0.96 0.00 0.00
-Average 0.98 0.96 0.00 0.98 0.00 0.98 0.92 0.94 0.00 0.93 0.00 0.39 0.97 0.97 0.38 0.97 0.00 0.46
-Table 3: Task-wise LIBER0 and LIBERO-PRO performance of three VLA models on the libero-spatial benchmark.
-Action notation: $Pick(src)$ = pick up object $bowl_{black}$ from location srcsrc and place $bowl_{black}$ onto/into target dstdst.
-Task (Symbolic Form) OpenVLA Pi0 Pi0.5
-Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env
-$Pick(between(plate,ramekin),plate)$ 1.00 0.95 0.00 0.88 0.00 0.02 1.00 0.97 0.00 0.94 0.00 1.00 1.00 0.96 0.02 0.98 0.00 0.66
-Pick(table_center,plate)Pick(table\_center,plate) 1.00 0.98 0.00 1.00 0.00 1.00 1.00 1.00 0.00 0.98 0.00 0.98 1.00 1.00 0.00 0.98 0.02 0.36
-$Pick(drawer_{top}(cabinet_{wood}),plate)$ 0.96 0.96 0.00 0.94 0.00 0.94 0.98 0.99 0.00 0.94 0.00 0.54 1.00 1.00 0.00 1.00 0.00 0.64
-Pick(next_to(cookie_box),plate)Pick(next\_to(cookie\_box),plate) 1.00 1.00 0.00 1.00 0.00 1.00 1.00 0.94 0.00 1.00 0.00 1.00 1.00 0.98 0.00 1.00 0.02 0.86
-Pick(next_to(plate),plate)Pick(next\_to(plate),plate) 1.00 0.95 0.00 0.94 0.00 0.92 0.96 1.00 0.00 0.94 0.00 0.84 0.90 0.92 0.00 0.90 0.00 0.22
-Pick(next_to(ramekin),plate)Pick(next\_to(ramekin),plate) 1.00 0.99 0.00 1.00 0.00 0.98 1.00 0.80 0.00 1.00 0.00 0.48 1.00 0.96 0.12 1.00 0.02 0.00
-Pick(on(cookie_box),plate)Pick(on(cookie\_box),plate) 1.00 0.99 0.00 0.98 0.00 1.00 0.98 0.92 0.00 0.98 0.00 0.96 1.00 1.00 0.00 0.98 0.00 1.00
-$Pick(on(ramekin),plate)$ 0.92 0.95 0.00 0.96 0.00 1.00 0.90 1.00 0.00 0.96 0.00 0.10 1.00 1.00 0.98 1.00 0.02 0.00
-$Pick(on(stove),plate)$ 1.00 0.99 0.00 1.00 0.00 1.00 0.96 0.99 0.00 0.96 0.00 0.04 0.96 0.96 0.02 0.94 0.00 0.82
-$Pick(on(cabinet_{wood}),plate)$ 0.94 0.99 0.00 1.00 0.00 1.00 0.94 0.86 0.00 1.00 0.00 0.80 0.90 0.88 0.90 0.90 0.00 0.02
-Average 0.98 0.97 0.00 0.97 0.00 0.89 0.97 0.95 0.00 0.97 0.00 0.60 0.98 0.97 0.20 0.97 0.01 0.46
-Table 4: Task-wise LIBERO and LIBERO-PRO performance of three VLA models on the libero-10 benchmark.
-Action notation: $TurnOn(obj)$ = activate object objobj; $Put(obj,loc)$ = place object objobj onto/into location locloc; $Place(obj,loc)$ = place picked object objobj into/onto target locloc; $(obj).close$ = close container objobj.
-Task (Symbolic Form) OpenVLA Pi0 Pi0.5
-Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env
-$TurnOn(stove)\wedge Put(moka\_pot,stove)$ 1.00 0.95 0.00 1.00 0.00 0.00 0.74 0.67 0.00 0.74 0.00 0.00 0.92 0.90 0.18 0.90 0.00 0.94
-$Put(bowl_{black},drawer_{bottom}(cabinet)).close$ 0.78 0.98 0.00 1.00 0.00 0.98 0.94 0.93 0.00 0.94 0.00 0.78 0.98 0.96 0.58 0.98 0.02 0.86
-$Put(mug_{yellow\_white},microwave).close$ 1.00 0.98 0.00 0.94 0.00 0.96 0.88 0.74 0.00 0.88 0.00 0.00 0.98 0.98 0.00 1.00 0.00 0.46
-$Put(\{moka\_pot_{1},moka\_pot_{2}\},stove)$ 0.92 0.64 0.00 0.92 0.00 0.78 0.22 0.28 0.00 0.22 0.00 0.00 0.98 0.98 0.00 0.98 0.02 0.26
-$Put(\{alphabet\_soup,cream\_cheese\},basket)$ 0.96 0.98 0.00 1.00 0.00 1.00 1.00 0.96 0.00 1.00 0.00 0.16 1.00 0.98 0.00 0.98 0.00 0.52
-$Put(\{alphabet\_soup,tomato\_sauce\},basket)$ 0.98 0.95 0.00 0.98 0.00 0.96 0.88 0.79 0.00 0.88 0.00 0.06 0.94 0.94 0.02 0.94 0.02 0.70
-$Put(\{cream\_cheese,butter\},basket)$ 0.96 0.96 0.00 0.98 0.00 0.98 0.98 0.98 0.00 0.98 0.00 0.96 0.96 0.94 0.00 0.98 0.00 0.00
-$Put((mug_{white},plate_{left})|(mug_{yellow\_white},plate_{right}))$ 0.84 0.85 0.00 0.92 0.00 0.98 0.86 0.72 0.00 0.86 0.00 0.00 1.00 0.98 0.00 1.00 0.02 0.22
-$Put((mug_{white},plate)|(pudding_{choco},right\_of(plate)))$ 0.94 0.77 0.00 0.90 0.00 0.86 0.76 0.82 0.00 0.74 0.00 0.00 0.62 0.58 0.00 0.60 0.00 0.04
-$Place(book,compartment_{back}(caddy))$ 1.00 1.00 0.00 0.98 0.00 1.00 0.94 0.98 0.00 0.94 0.00 0.76 0.94 0.94 0.00 0.94 0.00 0.56
-Average 0.93 0.81 0.00 0.96 0.00 0.85 0.82 0.79 0.00 0.82 0.00 0.27 0.93 0.92 0.08 0.93 0.01 0.46
-Table 5: Task-wise LIBERO and LIBERO-PRO performance of three VLA models on the libero-object benchmark.
-Action notation: $Place(obj,loc)$ = pick up object objobj and place objobj into/onto target locloc.
-Task (Symbolic Form) OpenVLA Pi0 Pi0.5
-Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env Ori Obj Pos Sem Task Env
-Place(alphabet_soup,basket)Place(alphabet\_soup,basket) 1.00 0.97 0.00 0.96 0.00 0.00 0.98 0.98 0.00 0.88 0.00 0.06 0.96 0.96 0.00 0.94 0.00 0.68
-Place(bbq_sauce,basket)Place(bbq\_sauce,basket) 0.98 0.90 0.00 0.94 0.00 0.00 0.98 1.00 0.00 1.00 0.00 0.46 1.00 0.98 1.00 1.00 0.02 0.96
-$Place(butter,basket)$ 1.00 1.00 0.00 1.00 0.00 0.00 1.00 0.98 0.00 1.00 0.00 0.00 0.96 0.98 0.54 0.98 0.00 0.00
-Place(chocolate_pudding,basket)Place(chocolate\_pudding,basket) 0.98 1.00 0.00 0.98 0.00 0.00 0.98 0.96 0.00 1.00 0.00 0.18 1.00 0.96 0.00 0.94 0.02 0.82
-Place(cream_cheese,basket)Place(cream\_cheese,basket) 1.00 1.00 0.00 1.00 0.00 0.00 1.00 1.00 0.10 1.00 0.00 0.76 0.98 1.00 0.00 0.94 0.00 0.88
-$Place(ketchup,basket)$ 1.00 1.00 0.00 1.00 0.00 0.00 0.96 0.84 0.00 0.92 0.00 0.24 0.98 0.96 0.20 0.96 0.02 0.92
-$Place(milk,basket)$ 1.00 0.98 0.00 0.96 0.00 0.00 1.00 0.84 0.00 1.00 0.00 0.16 1.00 1.00 0.00 0.98 0.00 0.80
-Place(orange_juice,basket)Place(orange\_juice,basket) 1.00 1.00 0.00 1.00 0.00 0.00 0.92 1.00 0.00 1.00 0.00 0.62 0.98 1.00 0.00 0.96 0.02 0.28
-Place(salad_dressing,basket)Place(salad\_dressing,basket) 0.94 1.00 0.00 1.00 0.00 0.00 0.98 0.98 0.10 1.00 0.00 0.00 1.00 0.98 0.00 0.98 0.00 1.00
-Place(tomato_sauce,basket)Place(tomato\_sauce,basket) 0.96 1.00 0.00 1.00 0.00 0.00 0.98 0.82 0.00 0.92 0.00 0.48 0.98 0.98 0.00 0.96 0.00 0.92
-Average 0.99 0.98 0.00 0.98 0.00 0.00 0.98 0.94 0.00 0.90 0.00 0.29 0.98 0.98 0.17 0.96 0.01 0.73
-5.3 Does the Model Generalize to New Objects?
-(a) "salad dressing" in the training set.
-(b) "salad dressing" not in the training set.
-(c) Replace "salad dressing" with "alphabet soup".
-Figure 4: Pick up salad dressing and place it in basket.
-To investigate whether the VLA model exhibits true object-level generalization, we conducted tests by altering the manipulated object in the task “pick up the salad dressing and place it in the basket”.
-As shown in Figure 4, replacing the original salad dressing with an unseen but visually similar variant from outside the training set (second row) did not affect the model’s performance, suggesting a limited degree of generalization to perturbed object instances.
-However, when substituting salad dressing with an unrelated object (e.g., alphabet soup, third row), the model still executed the same action trajectory, attempting to pick up the new object despite its mismatch with the instruction.
-This behavior suggests that the model’s apparent generalization is largely driven by rote recall of action trajectories from training rather than genuine grounding of object identity.
-Moreover, its failure on semantically irrelevant objects highlights the inability to correctly align linguistic references with object semantics.
-Taken together, these observations indicate that the model’s performance reflects superficial pattern memorization rather than true object-level reasoning.
-5.4 Does the Model Generalize to Varied Instructions?
-(a) Original task.
-(b) Semantic perturbation: Grab salad dressing and set it in basket.
-(c) Meaningless task: fdsgfdsgsd or xxxxxx.
-Figure 5: Pick up salad dressing and place it in basket.
-We evaluated whether the VLA model robustly follows natural language instructions.
-As shown in Figure 5, the model successfully executed the original task (“pick up the salad dressing and place it in the basket”) and also completed the task when the instruction was reformulated with synonymous expressions (“grab the salad dressing and set it in the basket”).
-Although this outcome may appear to suggest robustness to linguistic variation, when the instruction was replaced with meaningless input (e.g., “xxx”), the model produced the same action sequence, again retrieving and executing the pattern of placing the salad dressing into the basket.
-This observation indicates that the apparent robustness does not arise from genuine semantic comprehension.
-Rather, the model relies predominantly on visual context and memorized action trajectories from training, while disregarding the linguistic content of the input.
-Consequently, the model’s consistent behavior across both paraphrased and nonsensical instructions should not be regarded as successful generalization to semantic perturbations, but instead as evidence of its inability to ground or interpret instructions, revealing an overdependence on rote pattern execution.
-5.5 How Sensitive is the Model to Object Placement?
-Figure 6: Success rates of OpenVLA, pi0, and pi0.5 under object position perturbations.
-Although all three models achieve around 98% accuracy under the standard LIBERO, our perturbation experiments reveal significant robustness gaps that remain hidden in the original evaluation.
-As shown in Figure 6, the models are highly sensitive to the initial position of objects: for both OpenVLA and pi0, performance collapses once displacement exceeds 0.2 units, dropping to zero success, while pi0.5—benefiting from a stronger architecture and broader training data—exhibits improved resilience and sustains non-trivial accuracy up to 0.4 units before also collapsing to zero.
-These results indicate that high scores do not guarantee reliable performance in practice, as even small variations such as placing a cup in a slightly different position can cause complete task failure, underscoring the necessity of LIBERO-PRO to systematically reveal such robustness gaps and provide a more faithful comparison across models.
-5.6 Is the Model Truly Understanding Tasks?
-Figure 7: The model successfully executed two independent instructions, but failed to execute the combined instruction.
-To assess whether the model demonstrates genuine task comprehension or merely memorizes action trajectories, we designed a sequence of three progressively structured tasks (as shown in Figure 7).
-In the first task, the agent is instructed to pick up the bowl and place it on the stove, and the model executes this sequence successfully, indicating its ability to reproduce simple manipulation behaviors.
-In the second task, the instruction is to turn on the stove, and again the model succeeds, suggesting competence in carrying out isolated atomic actions.
-However, in the third task, which composes these two operations into a longer sequence—first picking up the bowl, then placing it on the stove, and finally turning the stove on—the model consistently fails.
-Rather than demonstrating adaptive reasoning over the composed instruction, the model replays spurious trajectories resembling those seen during training, leading to unsuccessful execution.
-This discrepancy reveals a critical limitation: the model does not truly understand either the semantic goal of the task or the causal relationships between subtasks, but instead exhibits rote reproduction of memorized patterns from the training distribution, and the failure on even modestly composed instructions underscores the fragility of current approaches and highlights the need for evaluation protocols that probe beyond single-step or template-like behaviors.
+6 Related Work
+Vision-Language-Action Models.
+Vision-Language-Action (VLA) models have rapidly emerged as a central paradigm for embodied AI, aiming to unify perception, natural language understanding, and action execution within a single framework.
+Recent years have witnessed an explosion of approaches, ranging from large-scale pre-trained models that leverage multimodal corpora to task-specific architectures tailored for robotic manipulation and navigation[14, 10, 2].
+While these efforts have significantly advanced the field, they also underscore the urgent need for standardized evaluation benchmarks[13, 15, 7].
+Without such benchmarks, it becomes difficult to conduct fair and reproducible comparisons across methods, hindering the cumulative progress of VLA research.
+Benchmarks for VLA Evaluation.
+A number of benchmarks have been developed to enable systematic evaluation of VLA models, including RLBench [7], CALVIN [15], RoboCasa [16], and BridgeData [19].
+These benchmarks provide valuable testbeds for task execution under vision and language guidance, yet their adoption has remained relatively limited within the broader community.
+In contrast, LIBERO [13] has quickly emerged as the dominant benchmark and is now the most widely adopted protocol for evaluating VLA models.
+Its standardized task suite and unified reporting metrics have made it the de facto standard for performance comparison, to the extent that virtually all recent VLA studies report results on LIBERO.
+As such, LIBERO not only serves as the “common currency” of evaluation in this field, but also exerts a strong influence on research directions and claims of progress.
+Limitations of Current Evaluation Practices.
+Several recent studies have noted that the high scores reported on LIBERO often fail to translate into reliable performance, even on tasks with only minor variations.
+This has raised serious concerns about the validity of its evaluation protocol[21, 3, 11].
+In response, some researchers have explored stronger algorithms or proposed stress tests [4, 5, 18, 9, 20]to encourage more robust VLA systems.
+However, the field still lacks a systematic and standardized framework for robust evaluation under LIBERO.
+The majority of works continue to rely on the original, flawed setup, reporting numbers that may not reflect genuine capability.
+To address this gap, we propose LIBERO-PRO, a plug-and-play extension of LIBERO designed to foster fair comparison across models and to advance research on building VLA systems that can withstand reasonable and practically necessary perturbations.
 7 Conclusion